OpenStudy (anonymous):

Determine \[k\] such that \[F\] is a distribution function: \[F(x,y)=kxy(2x+3y)\], where \[x+y le 1, x > 0, y > 0\]

OpenStudy (anonymous):

That is, \[F(x,y)=kxy(2x+3y), x + y \le 1, x>0, y>0\]

OpenStudy (zarkon):

Solve \[\int\limits_{0}^{1}\int\limits_{0}^{1-x}kxy(2x+3y)dydx=1\] for k the limit on the inner integral is 1-x (LaTeX need to be fixed)

myininaya (myininaya):

it looks fine to me

OpenStudy (anonymous):

Will try that. The limit was confusing me on that.

OpenStudy (zarkon):

you can see the 1-x ?

myininaya (myininaya):

yes

OpenStudy (zarkon):

hmmm...I cannot

myininaya (myininaya):

satellite you can see it right?

OpenStudy (anonymous):

Looks fine to me too

myininaya (myininaya):

refresh your page maybe zarkon? but gj on setting the problem up i knew it something to do with 1

OpenStudy (zarkon):

I'm using firefox (6.0beta) though I also had this issue with firefox v5

OpenStudy (anonymous):

i can see it. total has to be one good job!. i think inner integral is \[kx(x-1)^2\] but my integration stinks.

OpenStudy (zarkon):

I refresh and I still have the same issue

OpenStudy (zarkon):

you are correct satellite

OpenStudy (anonymous):

Not sure if any of you are still around, but I have a follow-up to this question. Took a break for lunch, and now I think I misunderstood the original question. The method given above would work if this were a probability density function, but the function given above is a cumulative distribution function. To use the method described above, would I first need to find the density function?

OpenStudy (zarkon):

that's what I would do

OpenStudy (anonymous):

Would the density function be\[k(4x+6y)\] I'm second-guessing myself here...

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

Ok. Thank you.